# An Information Theoretic Perspective on Conformal Prediction

### v1

@@ F.1 Centralized Setting
Table 6: Hyperparameter Search MNIST MLP
Bound Optmized for batch size lr temperature steepness THR APS Test Acc.
CE THR 100 0.01 - - $2.15_{\pm 0.21}$ $2.40_{\pm 0.13}$ 0.93
APS 500 0.05 - - $2.15_{\pm 0.21}$ $2.41_{\pm 0.12}$ 0.93
Fano THR 100 0.005 0.1 100 $2.08_{\pm 0.15}$ $2.11_{\pm 0.13}$ 0.91
APS 100 0.01 0.5 100 $2.05_{\pm 0.14}$ $2.08_{\pm 0.13}$ 0.91
MB Fano THR 1000 0.005 0.1 100 $2.15_{\pm 0.11}$ $2.83_{\pm 0.11}$ 0.91
APS 100 0.005 0.1 10 $2.31_{\pm 0.22}$ $2.37_{\pm 0.24}$ 0.92
DPI THR 1000 0.005 0.01 100 $2.18_{\pm 0.13}$ $2.81_{\pm 0.12}$ 0.92
APS 500 0.05 0.1 100 $2.21_{\pm 0.19}$ $2.52_{\pm 0.12}$ 0.93
ConfTr THR 500 0.05 0.1 100 $2.10_{\pm 0.16}$ $9.79_{\pm 0.01}$ 0.90
APS 500 0.01 1.0 100 $2.07_{\pm 0.15}$ $2.10_{\pm 0.16}$ 0.90
ConfTr-class THR 500 0.005 1.0 10 $2.05_{\pm 0.15}$ $2.10_{\pm 0.13}$ 0.90
APS 500 0.01 0.1 100 $2.08_{\pm 0.11}$ $2.08_{\pm 0.14}$ 0.90
Table 7: Hyperparameter Search Fashion-MNIST MLP
Bound Optmized for batch size lr temperature steepness THR APS Test Acc.
CE THR 1000 0.005 - - $2.18_{\pm 0.15}$ $2.48_{\pm 0.10}$ 0.87
APS 100 0.05 - - $2.04_{\pm 0.16}$ $2.24_{\pm 0.21}$ 0.89
Fano THR 100 0.05 0.5 100.0 $1.73_{\pm 0.03}$ $2.11_{\pm 0.08}$ 0.88
APS 100 0.01 1.0 100.0 $1.74_{\pm 0.05}$ $1.87_{\pm 0.06}$ 0.88
DPI THR 100 0.05 0.01 10.0 $1.75_{\pm 0.08}$ $2.15_{\pm 0.07}$ 0.90
APS 100 0.05 0.01 100.0 $1.68_{\pm 0.06}$ $2.06_{\pm 0.08}$ 0.89
MB Fano THR 100 0.05 0.5 10.0 $1.77_{\pm 0.06}$ $2.64_{\pm 0.12}$ 0.89
APS 100 0.05 1.0 100.0 $1.77_{\pm 0.09}$ $2.25_{\pm 0.12}$ 0.90
ConfTr THR 100 0.01 1.0 1.0 $1.69_{\pm 0.06}$ $1.87_{\pm 0.08}$ 0.89
APS 100 0.01 1.0 1.0 $1.69_{\pm 0.06}$ $1.87_{\pm 0.08}$ 0.89
ConfTr-class THR 100 0.01 1.0 10.0 $1.73_{\pm 0.06}$ $9.07_{\pm 2.37}$ 0.88
APS 500 0.005 1.0 100.0 $1.71_{\pm 0.03}$ $1.73_{\pm 0.05}$ 0.88
Table 8: Hyperparameter Search EMNIST MLP
Bound Optmized for batch size lr temperature steepness THR APS Test Acc.
CE THR 100 0.01 - - $2.01_{\pm 0.06}$ $3.25_{\pm 0.15}$ 0.86
APS 100 0.005 - - $2.00_{\pm 0.07}$ $3.27_{\pm 0.16}$ 0.86
Fano THR 100 0.01 0.1 100 $2.02_{\pm 0.08}$ $9.69_{\pm 0.57}$ 0.84
APS 100 0.01 1.0 100 $1.99_{\pm 0.06}$ $2.68_{\pm 0.10}$ 0.82
DPI THR 100 0.05 0.5 100 $1.98_{\pm 0.06}$ $3.85_{\pm 0.21}$ 0.86
APS 100 0.05 0.1 100 $1.99_{\pm 0.06}$ $4.13_{\pm 0.25}$ 0.86
MB Fano THR 100 0.01 1.0 100 $2.86_{\pm 0.22}$ $3.67_{\pm 0.15}$ 0.86
APS 100 0.01 0.1 1 $2.86_{\pm 0.22}$ $3.67_{\pm 0.15}$ 0.86
ConfTr THR 100 0.01 0.1 100 $1.95_{\pm 0.05}$ $5.04_{\pm 0.26}$ 0.85
APS 100 0.005 1.0 100 $1.96_{\pm 0.06}$ $2.32_{\pm 0.06}$ 0.83
ConfTr-class THR 100 0.005 0.1 100 $1.94_{\pm 0.05}$ $5.51_{\pm 0.41}$ 0.86
APS 100 0.05 1.0 100 $1.96_{\pm 0.05}$ $2.29_{\pm 0.06}$ 0.84
Table 9: Hyperparameter Search CIFAR10
Bound Optmized for batch size lr temperature steepness THR APS Test Acc.
CE THR 100 0.05 - - $1.81_{\pm 0.20}$ $2.34_{\pm 0.36}$ 0.93
APS 100 0.05 - - $1.83_{\pm 0.10}$ $2.41_{\pm 0.25}$ 0.93
Fano THR 100 0.01 1.0 1 $2.00_{\pm 0.10}$ $2.28_{\pm 0.08}$ 0.89
APS 100 0.01 1.0 1 $2.02_{\pm 0.06}$ $2.35_{\pm 0.10}$ 0.89
MB Fano THR 100 0.05 0.5 100 $1.70_{\pm 0.04}$ $2.35_{\pm 0.09}$ 0.92
APS 100 0.01 1.0 10 $1.74_{\pm 0.07}$ $2.01_{\pm 0.08}$ 0.92
DPI THR 100 0.05 0.01 100 $1.71_{\pm 0.07}$ $1.93_{\pm 0.12}$ 0.92
APS 100 0.005 0.01 10 $1.87_{\pm 0.12}$ $1.99_{\pm 0.10}$ 0.90
ConfTr THR 100 0.05 0.5 10 $9.77_{\pm 0.02}$ $9.98_{\pm 0.05}$ 0.08
APS 1000 0.005 0.1 1 $10.00_{\pm 0.00}$ $9.98_{\pm 0.00}$ 0.1
ConfTr-class THR 100 0.01 0.5 10 $2.29_{\pm 0.07}$ $2.27_{\pm 0.06}$ 0.85
APS 100 0.01 0.5 10 $2.14_{\pm 0.09}$ $2.24_{\pm 0.09}$ 0.86
Table 10: Hyperparameter Search CIFAR100
Bound Optmized for batch size lr temperature steepness THR APS Test Acc.
CE THR 100 0.05 - - $21.11_{\pm 1.16}$ $28.11_{\pm 2.38}$ 0.71
APS 100 0.05 - - $21.11_{\pm 1.16}$ $28.11_{\pm 2.38}$ 0.71
Fano THR 100 0.005 0.5 1 $36.60_{\pm 2.63}$ $44.39_{\pm 1.89}$ 0.50
APS 500 0.05 1 1 $34.51_{\pm 2.33}$ $37.01_{\pm 2.07}$ 0.55
MB Fano THR 100 0.05 0.5 100 $18.28_{\pm 1.63}$ $22.71_{\pm 0.93}$ 0.68
APS 100 0.05 1 10 $16.60_{\pm 1.04}$ $21.27_{\pm 1.38}$ 0.68
DPI THR 100 0.05 1.0 1 $16.03_{\pm 1.02}$ $19.07_{\pm 0.83}$ 0.71
APS 100 0.05 1 10 $15.54_{\pm 1.20}$ $16.86_{\pm 1.19}$ 0.69
ConfTr THR 100 0.005 1.0 1 $29.84_{\pm 2.37}$ $30.41_{\pm 2.69}$ 0.52
APS 100 0.01 0.5 1 $33.76_{\pm 1.97}$ $42.44_{\pm 2.05}$ 0.54
ConfTr-class THR 100 0.01 1.0 1 $31.25_{\pm 1.41}$ $49.62_{\pm 14.22}$ 0.43
APS 100 0.01 1 10 $32.57_{\pm 1.89}$ $32.65_{\pm 1.88}$ 0.37
F.2 Federated Setting
In the federated learning setting, we run conformal training exactly in the same fashion, but including the additional $Q_{Z|X}$ term in (10) to get the proper distributed bound that can be optimized locally and independently by each device. We optimize the conformal training objective with SGD for one epoch in each device, and then communicate the resulting “personalized” model to the server, which aggregates the model parameters of each device via federated averaging McMahan et al. (2017). After aggregation, the global model thus computed is communicated by to the client and the process restarts. We do 5K such communication rounds for EMNIST, and 10K for CIFAR10 and CIFAR100. As described in the main text, we divide the data among 100, 500, and 1K clients for CIFAR10, CIFAR100 and EMNIST, respectively. We assign data points to devices imposing a distribution-based label imbalance Li et al. (2022), where we sample a marginal label distribution for each device from a Dirichlet distribution. We use $\text{Dir}(1.0)$ for all experiments, but also study the effect of $\text{Dir}(0.5)$ and $\text{Dir}(0.1)$ on CIFAR10, as shown in Tables 11 and 12. The remainder of the experimental setup is similar to that used for the centralized setting, with the same architectures, data augmentation strategies, and optimal hyperparameters reported in Tables 6, 7, 8, 9. In Tables 11 and 12, as well as 13 and 14, we report inefficiency results for two different settings.

### v2

@@ G.1 Centralized Setting
Table 5: Hyperparameter Search for MNIST.
Bound Optimized for batch size lr temperature steepness THR APS Test Acc.
CE THR 100 0.01 - - $2.29_{\pm 0.18}$ $2.50_{\pm 0.08}$ 0.93
APS 500 0.05 - - $2.28_{\pm 0.19}$ $2.50_{\pm 0.08}$ 0.93
ConfTr THR 500 0.05 0.1 100 $6.28_{\pm 0.71}$ $9.81_{\pm 0.06}$ 0.90
APS 500 0.01 1.0 100 $2.08_{\pm 0.10}$ $2.10_{\pm 0.07}$ 0.90
ConfTr-class THR 500 0.005 1.0 10 $2.09_{\pm 0.11}$ $2.15_{\pm 0.12}$ 0.91
APS 500 0.01 0.1 100 $2.12_{\pm 0.11}$ $2.13_{\pm 0.13}$ 0.90
Fano THR 100 0.005 0.1 100 $2.09_{\pm 0.12}$ $2.12_{\pm 0.09}$ 0.91
APS 100 0.01 0.5 100 $2.11_{\pm 0.08}$ $2.12_{\pm 0.08}$ 0.91
MB Fano THR 1000 0.005 0.1 100 $2.24_{\pm 0.12}$ $2.96_{\pm 0.11}$ 0.91
APS 100 0.005 0.1 10 $2.43_{\pm 0.22}$ $2.49_{\pm 0.19}$ 0.92
DPI THR 1000 0.005 0.01 100 $2.24_{\pm 0.17}$ $2.87_{\pm 0.12}$ 0.92
APS 500 0.05 0.1 100 $2.29_{\pm 0.17}$ $2.64_{\pm 0.07}$ 0.92
Table 6: Hyperparameter Search for Fashion-MNIST.
Bound Optimized for batch size lr temperature steepness THR APS Test Acc.
CE THR 1000 0.005 - - $2.39_{\pm 0.13}$ $2.74_{\pm 0.20}$ 0.87
APS 100 0.05 - - $2.18_{\pm 0.17}$ $2.41_{\pm 0.17}$ 0.89
ConfTr THR 100 0.01 1.0 1.0 $1.73_{\pm 0.06}$ $1.89_{\pm 0.09}$ 0.89
APS 100 0.01 1.0 1.0 $1.73_{\pm 0.06}$ $1.89_{\pm 0.09}$ 0.89
ConfTr-class THR 100 0.01 1.0 10.0 $5.11_{\pm 0.49}$ $8.28_{\pm 3.17}$ 0.88
APS 500 0.005 1.0 100.0 $1.79_{\pm 0.06}$ $1.79_{\pm 0.07}$ 0.88
Fano THR 100 0.05 0.5 100.0 $1.70_{\pm 0.05}$ $2.06_{\pm 0.08}$ 0.88
APS 100 0.01 1.0 100.0 $1.76_{\pm 0.04}$ $1.87_{\pm 0.05}$ 0.87
MB Fano THR 100 0.05 0.5 10.0 $1.80_{\pm 0.08}$ $2.71_{\pm 0.14}$ 0.89
APS 100 0.05 1.0 100.0 $1.84_{\pm 0.10}$ $2.25_{\pm 0.14}$ 0.89
DPI THR 100 0.05 0.01 10.0 $1.73_{\pm 0.07}$ $2.12_{\pm 0.08}$ 0.89
APS 100 0.05 0.01 100.0 $1.75_{\pm 0.05}$ $2.08_{\pm 0.06}$ 0.89
Table 7: Hyperparameter Search for EMNIST.
Bound Optimized for batch size lr temperature steepness THR APS Test Acc.
CE THR 100 0.01 - - $2.06_{\pm 0.11}$ $3.37_{\pm 0.15}$ 0.86
APS 100 0.005 - - $2.06_{\pm 0.10}$ $3.40_{\pm 0.18}$ 0.86
ConfTr THR 100 0.01 0.1 100 $1.99_{\pm 0.10}$ $4.94_{\pm 0.26}$ 0.85
APS 100 0.005 1.0 100 $1.99_{\pm 0.08}$ $2.36_{\pm 0.11}$ 0.83
ConfTr-class THR 100 0.005 0.1 100 $2.01_{\pm 0.09}$ $5.07_{\pm 0.24}$ 0.85
APS 100 0.05 1.0 100 $1.99_{\pm 0.10}$ $2.38_{\pm 0.11}$ 0.84
Fano THR 100 0.01 0.1 100 $2.10_{\pm 0.11}$ $8.27_{\pm 0.53}$ 0.84
APS 100 0.01 1.0 100 $2.05_{\pm 0.11}$ $2.75_{\pm 0.14}$ 0.82
MB Fano THR 100 0.01 1.0 100 $2.01_{\pm 0.11}$ $5.21_{\pm 0.26}$ 0.86
APS 100 0.01 0.1 1 $2.94_{\pm 0.17}$ $3.67_{\pm 0.13}$ 0.86
DPI THR 100 0.05 0.5 100 $1.98_{\pm 0.09}$ $3.86_{\pm 0.20}$ 0.86
APS 100 0.05 0.1 100 $2.04_{\pm 0.12}$ $4.07_{\pm 0.23}$ 0.86
Table 8: Hyperparameter Search for CIFAR10.
Bound Optimized for batch size lr temperature steepness THR APS Test Acc.
CE THR 100 0.05 - - $1.69_{\pm 0.11}$ $2.12_{\pm 0.21}$ 0.93
APS 100 0.05 - - $1.74_{\pm 0.07}$ $2.34_{\pm 0.22}$ 0.93
ConfTr THR 100 0.05 0.5 10 $9.90_{\pm 0.02}$ $10.00_{\pm 0.00}$ 0.10
APS 1000 0.005 0.1 1 $9.90_{\pm 0.01}$ $9.98_{\pm 0.00}$ 0.10
ConfTr-class THR 100 0.01 0.5 10 $2.16_{\pm 0.09}$ $2.19_{\pm 0.10}$ 0.86
APS 100 0.01 0.5 10 $2.13_{\pm 0.08}$ $2.18_{\pm 0.06}$ 0.86
Fano THR 100 0.01 1.0 1 $2.05_{\pm 0.05}$ $2.34_{\pm 0.09}$ 0.89
APS 100 0.01 1.0 1 $2.06_{\pm 0.10}$ $2.35_{\pm 0.10}$ 0.89
MB Fano THR 100 0.05 0.5 100 $1.66_{\pm 0.09}$ $2.40_{\pm 0.08}$ 0.92
APS 100 0.01 1.0 10 $1.69_{\pm 0.09}$ $1.89_{\pm 0.06}$ 0.91
DPI THR 100 0.05 0.01 100 $1.64_{\pm 0.07}$ $1.88_{\pm 0.05}$ 0.92
APS 100 0.005 0.01 10 $1.79_{\pm 0.12}$ $1.97_{\pm 0.08}$ 0.91
Table 9: Hyperparameter Search for CIFAR100.
Bound Optimized for batch size lr temperature steepness THR APS Test Acc.
CE THR 100 0.05 - - $19.70_{\pm 2.05}$ $26.02_{\pm 1.31}$ 0.72
APS 100 0.05 - - $19.70_{\pm 2.05}$ $26.02_{\pm 1.31}$ 0.72
ConfTr THR 100 0.005 1.0 1 $32.80_{\pm 2.75}$ $34.09_{\pm 2.54}$ 0.52
APS 100 0.01 0.5 1 $30.04_{\pm 1.36}$ $40.58_{\pm 1.23}$ 0.53
ConfTr-class THR 100 0.01 1.0 1 $66.48_{\pm 3.67}$ $54.30_{\pm 17.12}$ 0.43
APS 100 0.01 1 10 $33.32_{\pm 1.89}$ $32.91_{\pm 1.53}$ 0.37
Fano THR 100 0.005 0.5 1 $40.30_{\pm 1.10}$ $48.21_{\pm 1.26}$ 0.42
APS 500 0.05 1 1 $30.43_{\pm 1.61}$ $33.80_{\pm 0.93}$ 0.57
MB Fano THR 100 0.05 0.5 100 $14.61_{\pm 0.84}$ $21.69_{\pm 0.71}$ 0.70
APS 100 0.05 1 10 $16.36_{\pm 0.93}$ $21.68_{\pm 1.44}$ 0.68
DPI THR 100 0.05 1.0 1 $17.55_{\pm 1.33}$ $20.13_{\pm 0.78}$ 0.69
APS 100 0.05 1 10 $14.90_{\pm 0.80}$ $17.41_{\pm 0.62}$ 0.70
G.1.1 Generalization to other target coverage rates
In all experiments, we run conformal training with a target coverage rate of 99%, i.e., $\alpha=0.01$. It is then important to assess whether the performance of the resulting models deteriorates at different coverage rates, “overfitting” to the value of $\alpha$ used for training. In Table 10, we see how inefficiency varies with $\alpha$ at test time for models trained via conformal training with $\alpha=0.01$. In particular, we can contrast their performance against that of models trained via the CE loss, which is agnostic to the desired coverage rate. In all cases, our model-based Fano and DPI bound performs best with the THR and APS methods, respectively, proving conformal training is worthwhile even if the desired coverage rate might vary at test time. Still, as noticed in [59], there is a drop in performance in comparison to the CE loss for higher values of $\alpha$ at test time. This could be due to some degree of overfitting, but it could also be attributed to the conformal prediction problem becoming easier for lower coverage rates, thus reducing the gap between our bounds and the CE loss.
Table 10: Inefficiency results with varying $\alpha$ at test time. Average prediction set size on CIFAR100 for different $\alpha$ targets at test time, averaged across 10 random calib./test splits. All methods were only optimized for $\alpha{=}0.01$. The models used for THR and APS might not be the same according to the best hyperparameters found in Table 9. Lower is better.
Method $\alpha=0.01$ $\alpha=0.05$ $\alpha=0.1$
THR APS THR APS THR APS
CE $19.70_{\pm 2.05}$ $26.02_{\pm 1.31}$ $6.11_{\pm 0.34}$ $9.19_{\pm 0.34}$ $3.02_{\pm 0.10}$ $4.52_{\pm 0.12}$
ConfTr $32.80_{\pm 2.75}$ $40.58_{\pm 1.23}$ $12.25_{\pm 0.47}$ $21.60_{\pm 0.78}$ $7.13_{\pm 0.23}$ $14.58_{\pm 0.47}$
[…]
Fano $40.30_{\pm 1.10}$ $33.80_{\pm 0.93}$ $19.43_{\pm 0.80}$ $16.17_{\pm 0.49}$ $11.46_{\pm 0.58}$ $9.72_{\pm 0.25}$
MB Fano $\mathbf{14.61_{\pm 0.84}}$ $21.68_{\pm 1.44}$ $\mathbf{5.24_{\pm 0.13}}$ $9.25_{\pm 0.30}$ $\mathbf{2.88_{\pm 0.05}}$ $5.51_{\pm 0.14}$
DPI $17.55_{\pm 1.33}$ $\mathbf{17.41_{\pm 0.62}}$ $6.26_{\pm 0.20}$ $\mathbf{6.98_{\pm 0.32}}$ $3.33_{\pm 0.11}$ $\mathbf{4.08_{\pm 0.14}}$
G.1.2 Results with Regularized Adaptive Prediction Sets (RAPS)
We report additional results with a conformal method known as RAPS (regularized adaptive prediction sets) [4]. In a nutshell, RAPS works exactly like APS but adds a penalization term $\lambda_{\text{reg}}$ to the score of each label that would make the prediction set larger than $k_{\text{reg}}.$ In Table 11, we have results with RAPS hyperparameters set as $k_{\text{reg}}{=}1$ and $\lambda_{\text{reg}}{=}0.01$, where we can see a pattern of performance across conformal training objectives that is similar to what we previously observed for THR and APS. As expected, the regularization introduced by RAPS results in prediction sets that are smaller than those obtained via APS in almost all cases. Again, similarly to [59], in all of our experiments we perform conformal training by thresholding on log-probabilities, which we observed to work best. The results we report for thresholding on probabilities (THR), APS and now RAPS show that this translates well to other score functions as well. Nonetheless, the improvement we get via RAPS still does not offset the gains from conformal training. In most cases, RAPS applied to a model trained via CE produces less efficient prediction sets than APS applied to a model trained to optimize one of our bounds.
Table 11: Inefficiency restuls with RAPS. Average prediction set size with both APS and RAPS ($k_{\text{reg}}=1$ and $\lambda_{\text{reg}}=0.01$) methods for each conformal training objective and data set for a target $\alpha=0.01$. We report mean and standard deviation across 10 different calib./test splits and show in bold all values within one standard deviation of the best result. Both APS and RAPS results were computed with the same classifier. Lower is better.
Method MNIST F-MNIST EMNIST CIFAR 10 CIFAR 100
APS RAPS APS RAPS APS RAPS APS RAPS APS RAPS
CE $2.50_{\pm 0.08}$ $2.35_{\pm 0.16}$ $2.41_{\pm 0.17}$ $2.40_{\pm 0.17}$ $3.40_{\pm 0.18}$ $3.07_{\pm 0.16}$ $2.34_{\pm 0.22}$ $2.31_{\pm 0.22}$ $26.02_{\pm 1.31}$ $26.02_{\pm 1.31}$
ConfTr $\mathbf{2.10_{\pm 0.07}}$ $\mathbf{2.10_{\pm 0.07}}$ $\mathbf{1.89_{\pm 0.09}}$ $1.84_{\pm 0.09}$ $\mathbf{2.36_{\pm 0.11}}$ $\mathbf{2.19_{\pm 0.06}}$ $9.99_{\pm 0.00}$ $9.99_{\pm 0.00}$ $40.59_{\pm 1.23}$ $39.87_{\pm 1.02}$
[…]
Fano $\mathbf{2.12_{\pm 0.08}}$ $\mathbf{2.12_{\pm 0.08}}$ $\mathbf{1.87_{\pm 0.05}}$ $1.80_{\pm 0.04}$ $\mathbf{2.75_{\pm 0.14}}$ $2.50_{\pm 0.07}$ $2.35_{\pm 0.10}$ $2.20_{\pm 0.04}$ $33.80_{\pm 0.93}$ $33.75_{\pm 1.02}$
MB Fano $\mathbf{2.49_{\pm 0.19}}$ $\mathbf{2.48_{\pm 0.19}}$ $\mathbf{2.25_{\pm 0.14}}$ $2.07_{\pm 0.11}$ $3.67_{\pm 0.13}$ $3.67_{\pm 0.13}$ $\mathbf{1.89_{\pm 0.06}}$ $\mathbf{1.78_{\pm 0.07}}$ $\mathbf{21.68_{\pm 1.44}}$ $20.97_{\pm 1.30}$
DPI $\mathbf{2.64_{\pm 0.07}}$ $2.44_{\pm 0.12}$ $\mathbf{2.08_{\pm 0.06}}$ $1.95_{\pm 0.05}$ $\mathbf{4.07_{\pm 0.23}}$ $3.74_{\pm 0.17}$ $\mathbf{1.97_{\pm 0.08}}$ $\mathbf{1.87_{\pm 0.11}}$ $\mathbf{17.41_{\pm 0.62}}$ $\mathbf{17.25_{\pm 0.58}}$
We also study the impact of the two hyperparameter in RAPS, namely $\lambda_{reg}$ and $k_{reg}$ in the inefficiency of each of the conformal training methods. We concentrate our analysis on CIFAR100 and report results with varying $k_{\text{reg}}$ in Table 12 and varying $\lambda_{\text{reg}}$ in Table 13. Interestingly, RAPS was much more sensitive to variations in $k_{\text{reg}}$ than in $\lambda_{\text{reg}}$. Still, we observe models trained via our model-based Fano and DPI bounds produce more efficient prediction sets than the baselines in all cases.
Table 12: Inefficiency results for RAPS with varying $k_{\text{reg}}$. Average prediction set size for $\lambda_{\text{reg}}=0.01$ and varying $k_{\text{reg}}$ on CIFAR100 for a target $\alpha=0.01$. We report mean and standard deviation across 10 different calib./test splits and show in bold all values within one standard deviation of the best result. Both APS and RAPS results were computed with the same classifier. Lower is better.
Method $k_{\text{reg}}=0$ (APS) $k_{\text{reg}}=1$ $k_{\text{reg}}=3$ $k_{\text{reg}}=5$ $k_{\text{reg}}=10$ $k_{\text{reg}}=15$
CE $26.02_{\pm 1.31}$ $26.02_{\pm 1.31}$ $22.43_{\pm 1.88}$ $20.43_{\pm 1.70}$ $18.95_{\pm 1.18}$ $20.63_{\pm 0.89}$
ConfTr $40.59_{\pm 1.23}$ $39.87_{\pm 1.02}$ $38.73_{\pm 0.84}$ $38.31_{\pm 0.70}$ $38.06_{\pm 0.48}$ $38.06_{\pm 0.59}$
[…]
Fano $33.80_{\pm 0.93}$ $33.75_{\pm 1.02}$ $32.83_{\pm 1.60}$ $31.71_{\pm 1.55}$ $30.82_{\pm 1.72}$ $30.31_{\pm 1.71}$
MB Fano $21.68_{\pm 1.44}$ $20.97_{\pm 1.30}$ $19.32_{\pm 1.40}$ $19.04_{\pm 1.36}$ $\mathbf{17.35_{\pm 0.86}}$ $\mathbf{17.68_{\pm 0.94}}$
DPI $\mathbf{17.41_{\pm 0.62}}$ $\mathbf{17.25_{\pm 0.58}}$ $\mathbf{17.02_{\pm 0.75}}$ $\mathbf{16.40_{\pm 0.93}}$ $\mathbf{17.68_{\pm 0.82}}$ $19.49_{\pm 2.08}$
Table 13: Inefficiency results for RAPS with varying $\lambda_{\text{reg}}$. Average prediction set size for $k_{\text{reg}}=1$ and varying $\lambda_{\text{reg}}$ on CIFAR100 for a target $\alpha=0.01$. We report mean and standard deviation across 10 different calib./test splits and show in bold all values within one standard deviation of the best result. Both APS and RAPS results were computed with the same classifier. Lower is better.
Method $\lambda_{\text{reg}}=0$ (APS) $\lambda_{\text{reg}}=0.01$ $\lambda_{\text{reg}}=0.03$ $\lambda_{\text{reg}}=0.1$ $\lambda_{\text{reg}}=0.3$ $\lambda_{\text{reg}}=1.0$
CE $26.02_{\pm 1.31}$ $26.02_{\pm 1.31}$ $26.02_{\pm 1.31}$ $26.02_{\pm 1.31}$ $26.02_{\pm 1.31}$ $26.02_{\pm 1.31}$
ConfTr $40.59_{\pm 1.23}$ $39.87_{\pm 1.02}$ $39.17_{\pm 1.05}$ $39.09_{\pm 1.12}$ $39.09_{\pm 1.12}$ $39.09_{\pm 1.12}$
[…]
Fano $33.80_{\pm 0.93}$ $33.75_{\pm 1.02}$ $33.75_{\pm 1.02}$ $33.75_{\pm 1.02}$ $33.75_{\pm 1.02}$ $33.75_{\pm 1.02}$
MB Fano $21.68_{\pm 1.44}$ $20.97_{\pm 1.30}$ $20.96_{\pm 1.29}$ $20.96_{\pm 1.29}$ $20.96_{\pm 1.29}$ $20.96_{\pm 1.29}$
DPI $\mathbf{17.41_{\pm 0.62}}$ $\mathbf{17.25_{\pm 0.58}}$ $\mathbf{17.25_{\pm 0.58}}$ $\mathbf{17.25_{\pm 0.58}}$ $\mathbf{17.25_{\pm 0.58}}$ $\mathbf{17.25_{\pm 0.58}}$
G.2 Federated Setting
Regarding the federated learning setting, we start by providing extra results with APS in Table 14 to complement results with THR in Table 3 in the main text. Additionally, we run conformal training exactly in the same fashion, but including the additional $Q_{Z|X}$ term in (11) to get the proper distributed bound that can be optimized locally and independently by each device. We optimize the conformal training objective with SGD for one epoch in each device, and then communicate the resulting “personalized” model to the server, which aggregates the model parameters of each device via federated averaging [40]. After aggregation, the global model thus computed is communicated by to the client and the process restarts. We do 5K such communication rounds for EMNIST, and 10K for CIFAR10 and CIFAR100. As described in the main text, we divide the data among 100, 500, and 1K clients for CIFAR10, CIFAR100 and EMNIST, respectively. We assign data points to devices imposing a distribution-based label imbalance [31], where we sample a marginal label distribution for each device from a Dirichlet distribution. We use $\text{Dir}(1.0)$ for all experiments, but also study the effect of $\text{Dir}(0.5)$ and $\text{Dir}(0.1)$ on CIFAR10, as shown in Tables 15 and 16. The remainder of the experimental setup is similar to that used for the centralized setting, with the same architectures, data augmentation strategies, and optimal hyperparameters reported in Tables 5, 6, 7, 8. In Tables 15 and 16, as well as 17 and 18, we report inefficiency results for two different settings.
Table 14: Inefficiency results for conformal training in the federated setting with APS. We report the mean prediction set size ($\pm$ standard deviation) of the global federated model across 10 different calib./test splits for $\alpha=0.01$ and using APS. We use ${}_{+\text{SI}}$ to indicate the inclusion of side information. We show in bold all values within one standard deviation of the best result. Lower is better.
Method EMNIST CIFAR 10 CIFAR 100
APS APS${}_{+\text{SI}}$ APS APS${}_{+\text{SI}}$ APS APS${}_{+\text{SI}}$
CE $3.69_{\pm 0.03}$ $3.14_{\pm 0.04}$ $\mathbf{2.83_{\pm 0.07}}$ $2.43_{\pm 0.06}$ $64.73_{\pm 0.34}$ $62.67_{\pm 3.68}$
ConfTr $6.14_{\pm 0.04}$ $5.25_{\pm 0.04}$ $10.00_{\pm 0.00}$ $10.00_{\pm 0.00}$ $55.18_{\pm 2.10}$ $47.58_{\pm 1.48}$
[…]
Fano $3.12_{\pm 0.04}$ $2.72_{\pm 0.03}$ $\mathbf{2.73_{\pm 0.07}}$ $\mathbf{2.39_{\pm 0.06}}$ $\mathbf{46.95_{\pm 0.67}}$ $\mathbf{42.75_{\pm 0.91}}$
MB Fano $4.75_{\pm 0.03}$ $\mathbf{2.43_{\pm 0.01}}$ $\mathbf{2.79_{\pm 0.13}}$ $\mathbf{2.33_{\pm 0.05}}$ $50.72_{\pm 1.77}$ $45.72_{\pm 1.38}$
DPI $2.98_{\pm 0.03}$ $2.58_{\pm 0.02}$ $\mathbf{2.68_{\pm 0.15}}$ $\mathbf{2.22_{\pm 0.09}}$ $51.29_{\pm 1.07}$ $47.18_{\pm 1.27}$